OpenStudy (anonymous):

Write a quadratic equation having the given number as solutions. -5 and -1 The quadratic equation is ?=0

OpenStudy (anonymous):

y=(x+5)(x+1) ....y=x^2+6x+5

OpenStudy (anonymous):

(x+5)(x+1) x^2+6x+5